OpenStudy (anonymous):

I need to find the integral of x^3(x^4-4)^12. With U=((x^4)-4)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

((1/3)(u)^13)/13

OpenStudy (anonymous):

do substitution: u= x^4 - 4 so: u'=4x^3 so:\[=1/4 \int\limits_{}^{}u ^{12}du = 1/4 *1/13 *u ^{13}=1/52 *(x ^{4}-4)\]

OpenStudy (anonymous):

So, (x^4-4)^13/39

OpenStudy (anonymous):

oups... it should be \[1/52 * (x ^{4}-4)^{13}\]

OpenStudy (anonymous):

good call :))

OpenStudy (anonymous):

Yeah, its the , but sqme otherwise

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Lol sorry, my iPad hates this site so I can barely see it or type. Oops!